OpenStudy (anonymous):

Which function has the greatest starting value? A. x y 1 −3 2 −8 3 −13 B. y = 3x – 8 C. 8x + 4y = 16 D. A linear function that goes through point 2, 0 and point 0, 2

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

It's asking for the greatest y-intercept.

OpenStudy (igreen):

It says that D goes through the point (0, 2), which is the y-intercept..2.

OpenStudy (igreen):

y = 3x – 8 Do you know what the y-intercept of this is(the number after 'x')?

OpenStudy (anonymous):

8?

OpenStudy (igreen):

No, almost..it's -8, because 8 is being subtracted.

OpenStudy (anonymous):

o

OpenStudy (igreen):

8x + 4y = 16 Change it to slope-intercept form: Subtract 8x to both sides: 4y = -8x + 16 Divide 4 to both sides: y = -2x + 4 What's the y-intercept of this one?

OpenStudy (anonymous):

-2

OpenStudy (igreen):

Yep. And the y-intercept of A is 0. So: A is 0 B is -8 C is -2 D is 2 So which one is greatest?

OpenStudy (michele_laino):

I think, you have to evaluate the value of the functions of your exercise, at point x=1. Please, you have to write the form of linear function, in order to that, we can write: \[f(x)=ax+b\], and to determine the constants a, and b, using the coordinates of the two points that you have indicated in your problem

OpenStudy (anonymous):

sorry its D

OpenStudy (igreen):

Yep, you got it.